Please prove the following identityCos 4x=1+8cos^4x-8cos^2x

Grade:11

2 Answers

kumar
6 years ago
cos4X = cos(3x +x) = cos
kumar
6 years ago
cos4X = cos(3x +x) = cos3x. cosx – sin3x. sinx then use forumla for cos3x and sin3x and convert all terms in cosine format